OpenStudy (anonymous):

surface area of revolution y=sinpix on x=0 to x=1.

OpenStudy (anonymous):

\[y=sinpix \]

OpenStudy (anonymous):

f'x=(cospix)pi

OpenStudy (anonymous):

\[2\Pi \int\limits_{0}^{1}\sin Pix(\sqrt{1+(picospix)^2}\]

OpenStudy (anonymous):

I could be mistaken, but that last integral looks to me like what I would use for arc length. I thought for area we would integrate pi(r)^2, and in your case it looks like r is sin(pix).

OpenStudy (anonymous):

So if I am right, then you would have to square sin(pix) before integrating it. you'd have to check, but I suspect there is probably a helpful identity you could sub in for sin^2(pix)

OpenStudy (anonymous):

finding surface area of revolution involes arc length integral

OpenStudy (tkhunny):

Arc length in 2D is very much associated with Surface Area in 3D. The two integrals should have obvious similarities.

OpenStudy (anonymous):

looks like i missed the surface area part, my mistake.

OpenStudy (anonymous):

Still doesn't look right.

OpenStudy (anonymous):

What formula is being used?

OpenStudy (anonymous):

\[ 2\pi r \]Takes the radius and gives your circumference. Arc length isn't the radius.

OpenStudy (anonymous):

\[\int\limits_{a}^{b}2\pi(fx)\sqrt{1+[f'(x)]^2}\]

OpenStudy (anonymous):

this equation has been getting me the right answers so far

OpenStudy (tkhunny):

It is correct. It was also correct up at the top. Why did you give up on it?

OpenStudy (anonymous):

not entirely sure how to square picospix, then i have to add 1 to it, then i have to to take square root. how do i go about doing that

OpenStudy (anonymous):

is picospix squared = pi^2cos^2(pix)?

OpenStudy (anonymous):

yh, looks like it is, just chcked in calc

OpenStudy (tkhunny):

That is fine. Don't worry about the square root. It will not bother you. Try substituting \(u = \cos(\pi x)\)

OpenStudy (anonymous):

let u = cospix du = sinpixdx \[\int\limits_{a}^{b}du \sqrt{1+\pi^2u^2}\]

OpenStudy (anonymous):

whoops, forgot the chain rule taking the derivative

OpenStudy (tkhunny):

Missing a few parts, there. \(du = -\pi\sin(\pi x)\;dx\)

OpenStudy (anonymous):

let u=cospix du=-pisinpix -du/pi=sinpix \[-1/\pi \int\limits_{a}^{b}du \sqrt{1+\pi^2u^2}\]

OpenStudy (anonymous):

@tkhunny how about now? :p

OpenStudy (anonymous):

I'm definitely in over my head, but that looks a little bit now like it meets sqrt(a^2 + x^2) for a trig substitution

OpenStudy (tkhunny):

Well, it's still not pretty, but you should be able to wade through it to a complete resolution. Save a whole sheet of paper for the details. :-)

OpenStudy (anonymous):

it still looks completely un do able to me to l0l. and yeah, i think it might be a trig subsitution, but not entirely sure how i would go about do that

OpenStudy (tkhunny):

It was trig when we started. You might want to back up and try a different substitution. If all else fails, you can try Weierstrass - annoying, but it works.

OpenStudy (anonymous):

even if i let u=sinx, i just end up with another problem with a bunch of u's and pi's that i dont know what to do with lol. guess i'll just move to another problem for now

OpenStudy (anonymous):

thanks anyways though

OpenStudy (anonymous):

let u =sinpix* is what i meant

OpenStudy (tkhunny):

Yea, it's not pretty, but it can be done. Good work this far and good luck on the rest!